If F(u,v) = u-v/u+v, find F (1/u,1/v) + F (u,v).

1
Expert's answer
2022-06-08T13:40:04-0400
"F (1/u,1/v) + F (u,v)=\dfrac{1/u-1/v}{1/u+1/v}+\dfrac{u-v}{u+v}"

"=\dfrac{v-u}{u+v}+\dfrac{u-v}{u+v}=0, u\not=0, v\not=0, u\not=-v"